OpenStudy (anonymous):

Help, identify and state symmetries for the polar equation? I'll try graphing it on my own. Any help is great!

OpenStudy (anonymous):

r=9cos(5θ)

OpenStudy (kmeis002):

Polar equations that follow the form: \[ r = asin(n\theta) \text{ or } r = acos(n\theta) \] are generally called rose plots, namely because they look like flowers. The a is the largest distance of each petal, and the n is the number of petals with a small caveat. If n is odd, then the plot has n petals. If n is even, the the plot has 2n petals. Our plot should have a petal length of 9 and 5 petals. Since its cosine, the first petal should start on the x axis (sine plots would start on the y).

OpenStudy (anonymous):

thank you so much!